A chef is using a mixture of two brands of Italian dressing. The first brand contains
10%
vinegar, and the second brand contains
14%
vinegar. The chef wants to make
220
milliliters of a dressing that is
11%
vinegar. How much of each brand should she use?

Answer by stanbon(75887)

A chef is using a mixture of two brands of Italian dressing. The first brand contains 10% vinegar, and the second brand contains 14% vinegar. The chef wants to make 220 milliliters of a dressing that is 11% vinegar. How much of each brand should she use?
-------------
Equation:
vinegar + vinegar = vinegar
----
0.10x + 0.14(220-x) = 0.11*220
-----
10x + 14*220 - 14x = 11*220
-----
-4x = -3*220
x = (3/4)220
x = 3*55 = 165 ml (amt. of 10% needed)
220-165 = 55 ml (amt. of 14% needed)
------------
Cheers,
Stan H.
---------------